### 30 Mar 2026 v2

# The effect of interactions on elastic cavitation

Abstract
Cavitation refers to the sudden, unstable expansion of a defect or cavity within a material in response to applied loads, when the loads reach a critical threshold. It is widely recognized as a common failure nucleation mechanism in soft and biological materials. For an isolated cavity in the bulk of an incompressible neo-Hookean solid loaded by remote hydrostatic tension, the classical cavitation pressure is well established as $p_{\rm c,bulk}=2.5\mu$, where $\mu$ is the shear modulus. However, in realistic settings the cavitation threshold is influenced by additional factors, including proximity to interfaces and interactions with neighboring cavities. Interface interaction effects are particularly relevant in modern multi-material systems and additively manufactured structures, where defects frequently occur near material boundaries. Meanwhile, cavity-cavity interactions become important in materials exhibiting finite porosity, such as foams, porous solids, and phase-separating polymers.
In this study, we characterize the effect of cavity interaction on cavitation pressure for (i) a nearby rigid interface and (ii) a neighboring identical cavity. For cavities near a rigid interface, our analysis shows that the cavitation pressure increases as the initial cavity-interface distance decreases, starting from the bulk value $p_{\rm c,bulk}$ for a distant cavity and approaching the cavitation pressure value for a defect situated at an interface ($p_{\rm c,int}\approx 3.5\mu$) as the cavity approaches the interface boundary. In contrast, interacting cavities exhibit a non-monotonic dependence of the cavitation pressure on the initial inter-cavity distance $d$: the threshold approaches the bulk value $p_{\rm c,bulk}$ for distant cavities and reaches a maximum of $p_{\rm c}\approx 2.67\mu$ at $d\approx 3.8R$, where $R$ is the initial cavity radius. These results provide, for the first time, a quantitative estimate of cavitation pressure for these two interaction-effect extensions of the classical cavitation problem and will help guide modeling of cavitation-driven processes in porous, biological, and multi-material solids.
keywords:
Cavitation, Boundary effects, Interaction effects, Cavity growth, neo-Hookean material
1 Introduction
Cavitation is a mechanical instability in which small pre-existing voids or defects embedded in a solid expand suddenly and unstably once an applied load crosses a critical threshold. It is a critical failure nucleation mechanism in soft and biological materials [5] as well as metals [22]. The phenomenon was first systematically characterized in rubber-like materials by Gent and Lindley [13], who showed that spherical cavities embedded in an incompressible neo-Hookean solid grow without bound when the remotely applied hydrostatic tension reaches the critical value $p_{\rm c,bulk}=2.5\mu$, where $\mu$ is the shear modulus. This result was later reinterpreted by Ball [4] as a bifurcation of the equilibrium equations of nonlinear elasticity: a solid ball that is perfectly defect-free may spontaneously open a cavity at its center when the applied load is sufficiently large, yielding the same critical pressure as the defect-growth picture.
Since these foundational studies, cavitation has been investigated extensively across a wide range of material systems and loading scenarios. The role of material compressibility was examined by Horgan and Abeyaratne [19] for foam rubberlike (Blatz–Ko) materials, with further explicit radially symmetric solutions for compressible nonlinear elastic solids provided by Horgan [20]. Stuart [44], Sivaloganathan [43], and Meynard [36] addressed existence and uniqueness of cavitating equilibria for broader classes of compressible materials. Further extensions of radially symmetric cavitation to anisotropic materials were studied by Antman and Negrón-Marrero [3] and by Polignone and Horgan [39], while dynamic [38] and viscoelastic [11, 26] effects have also been investigated. A review of radially symmetric cavitation in nonlinear elastic solids can be found in [18].
Hou and Abeyaratne [21] extended cavitation analysis to non-symmetric loading by constructing an approximate cavitation surface in full principal-stress space, showing that the hydrostatic pressure criterion provides a useful but imperfect approximation when the loading deviates from purely volumetric tension. Lopez-Pamies [35] later derived a closed-form cavitation criterion for compressible, isotropic, hyperelastic solids under general plane-strain loading using a linear-comparison variational method. Lopez-Pamies and co-workers [33, 34] subsequently recast cavitation as a homogenization problem, recovering Ball’s classical result as a special case and extending it to arbitrary three-dimensional loading conditions, compressible and anisotropic materials, and general random distributions of pre-existing defects. In addition to vacuous cavities and defects, non-symmetric cavitation triggered by inclusions growing inside the material was studied by Bonavia and co-workers [6], where the resistance pressure faced by the growing inclusion saturates at a critical value with increasing volume of inclusion.
While these works provide increasingly general descriptions of cavitation, most remain focused on isolated growth of cavities in a perfect elastic medium. However, cavities rarely grow in isolation in biological and modern engineering materials. Soft composites contain inclusions and reinforcement particles, where growth of cavities near material interfaces affects cavitation and failure [14, 7]. Similarly in porous solids and biological tissues, and phase separating polymers, growing cavities can interact with each other which can affect cavitation and failure. In porous elastomers for example, Lopez-Pamies and Ponte Castañeda [32] showed that interactions among cavities encoded in the homogenized constitutive response can trigger macroscopic instabilities. Thus it is critical to study how cavitation is affected by the interaction of growing cavities with nearby interfaces and other growing cavities to robustly model and understand the behavior of porous, biological, and multi-material solids.
There are a limited number of cavitation studies that focus on interaction effects. Kang and coworkers [23] studied cavity expansion in inhomogeneous soft solids and demonstrated that proximity to a planar interface between two materials of differing moduli can significantly alter the cavitation response. In an earlier study, Henzel and co-workers [16] analyzed cavitation arising from a defect that exists at the interface between an elastic body and a bonded rigid substrate and established that the critical pressure rises to $p_{\rm c,int}\approx 3.5\mu$—a forty-percent increase over the bulk value attributable entirely to the constraint imposed by the rigid boundary. The framework of Lopez-Pamies and co-workers [33] can be used to account for inter-cavity interactions but only in an average homogenized sense through an assumed distribution of voids. There are no studies so far that have analyzed how the proximity of growing cavities to nearby interfaces and other growing cavities affects the cavitation limit. The current work addresses this gap.
In this manuscript, we study the effect of interactions on cavitation while neglecting surface tension and fracture physics (to isolate the effect of interactions). Specifically, we focus on two fundamental cases—the interaction of a growing cavity with a nearby rigid interface, and, separately, with a nearby identically growing cavity. Thus, we consider an incompressible neo-Hookean solid under remote hydrostatic tension loading and analyze two configurations: (i) a single spherical cavity with radius $R$ at a distance $d$ from a nearby planar rigid interface, and (ii) two identical spherical cavities of radii $R$ separated by distance $2d$. Our prior work [40, 41] characterized the energy landscape and configurational forces in the latter setting, but here we focus on the cavitation limit.
The paper is organized as follows. We begin by presenting the results of the study in §2, followed by a detailed description of the problem and solution setup used to obtain the results in §3. The computational framework is outlined in §4, and we end with concluding remarks in §5.
Figure 1: Axisymmetric schematic of the cavitation problems being studied: (a) cavity-interface interaction, and (b) cavity-cavity interaction. The cavities are under remote hydrostatic tension loading.
2 Results
Proximity to an interface or another cavity (Fig. 1) affects the cavitation pressure relative to its bulk value. To quantitatively describe these effects, we introduce the following notation. The dimensionless pressure is defined as $\overline{p}=p/\mu$, where $p$ is the remotely applied pressure and $\mu$ is the shear modulus of the material. The dimensionless volume ratio is defined as $\overline{V}=V/V_{0}$, where $V_{0}$ is the initial, undeformed cavity volume, and $V$ is its deformed volume. The critical value of pressure $p$ which causes unstable cavity expansion is the cavitation pressure $p_{\rm c}$, and $\overline{p}_{\rm c}=p_{\rm c}/\mu$ is the dimensionless cavitation pressure. As a reference value, the bulk cavitation pressure for a spherical cavity in an infinite, incompressible neo-Hookean medium is $\overline{p}_{\rm c,\rm bulk}=2.5$. The cavitation pressure for expansion of a defect located exactly at the interface between the elastic body and a bonded rigid substrate was established in a prior study [16] as $\overline{p}_{\rm c,int}\approx 3.5$, also called the interfacial cavitation pressure. The key results of this study are as follows:
Cavity near a rigid interface in a semi-infinite medium
Fig. 2 shows the pressure-volume relation for a sequence of distances (expressed in the dimensionless form $d/R$ where $R$ is the radius of the cavity and $d$ is the distance of its center from the rigid interface in the undeformed configuration) and Fig. 2 shows the corresponding post-processed cavitation pressures. The cavitation pressure increases monotonically as the undeformed cavity location gets closer to the interface, attaining $\overline{p}_{\rm c}\approx 3.44$ at d/R=1.1d/R=1.1—a significant increase from the bulk cavitation pressure (which is the asymptotic limit of the cavitation pressure for $d/R\to\infty$). As the distance between the cavity and the interface increases, the pressure-volume curve approaches that of the bulk cavitation case (dashed black curve in Fig. 2, defined in eq. (3)).
[…]
Figure 2: Cavitation near a rigid interface: (a) dimensionless pressure–volume curves for different dimensionless cavity–interface distances $d/R=\{1.1,3,6,20\}$. The dashed black curve shows analytical bulk cavitation case (eq. (3)); (b) dimensionless cavitation pressure as a function of $d/R$.
Two cavities in an infinite medium
The cavitation pressure varies non-monotonically with the undeformed inter-cavity half-distance $d$, as shown in Fig. 3. A peak is observed at $d/R\approx 3.8$ where $\overline{p}_{\rm c}\approx 2.67$, highlighting a maximal cavity interaction effect; for large $d/R$, the response approaches the single-cavity limit $\overline{p}_{\rm c}=2.5$ (dashed black curve in Fig. 3).
[…]
Figure 3: Cavitation of two interacting cavities: (a) dimensionless pressure-volume curves of two interacting cavities for different dimensionless inter-cavity half-distances $d/R=\{1.1,3,6,20\}$. The dashed black curve shows analytical bulk cavitation case (eq. (3)); (b) dimensionless cavitation pressure as a function of $d/R$.
The deformed shapes of the cavity have been plotted at several discrete $\overline{V}$ values for both the cavity-interface and cavity-cavity interaction problems in Figure 4.
Figure 4: Deformed cavity shapes in (top) cavity-interface interaction and (bottom) cavity-cavity interaction problems for different dimensionless volumes.
It is worth pointing out that the enhancement effect in cavitation pressures found in this study is meaningful for physical systems. While one may contend that, in theory, other “isolated” defects might cavitate earlier when the pressure reaches the (lower) bulk cavitation value, in real physical systems, surface energy effects impose severe penalties on the nucleation and growth of small defects [12, 29]. We assume that the cavities considered in our study are the most relevant size-critical defects in the system.
3 Problem Setup
3.1 Geometry and Kinematics
We investigate two problems in this work: i) a spherical cavity near a rigid interface in a semi-infinite medium, and ii) two equi-sized spherical cavities in an infinite medium. In both problems, the system is subjected to remote pressure (hydrostatic tension) loading [^1] The findings from this problem setup are also applicable to physical situations where the cavities are instead internally pressurized (such as from biological [27, 8, 42, 25, 45, 37] and other growth [24, 6] processes, thermo-chemical processes in energetic materials [9], or through injected fluid in cavity-based viscoelastic characterization techniques [10, 28]) since we can show equivalence in the displacement fields of the remote loading and internal loading problems [17].. These are both axisymmetric problems, whose 2D representations are shown in Figure 5, with $X_{2}$ being the axis of symmetry. In numerical simulations, the remote boundaries are implemented using large finite dimensions; the additional symmetry about the $X_{1}$ axis in the interacting-cavities problem allows us to reduce its geometry to the same as that of the cavity-interface interaction problem except that the bottom boundary is a symmetry plane instead of a rigid interface.
Figure 5: Axisymmetric schematic of the cavitation problem and its simulation model for the case of (a) cavity-interface interaction, and (b) cavity-cavity interaction. Both load-controlled and displacement-controlled representations are shown.
Approach
The cavitation problem is a load-controlled problem. However, since we do not know a priori the cavitation pressure to approach the limit in a controlled manner, the load-controlled problem is not very tractable. We employ a displacement-controlled approach to alleviate this issue. Given the infinite size of the domain, the displacements at the remote boundaries are expected to be uniform. The volume change of the cavity can then be directly related to the imposed uniform displacements since the material is incompressible. The equivalent displacement-control problem is shown in Figure 5.
The volume change at the top boundary must equal the volume change for cavity, which determines the displacement at the top boundary as
$u_{2}^{\mathrm{top}}=\frac{V-V_{0}}{\pi L^{2}},$ (1)
where $L$ is the length of the top edge of the simulation domain (see Fig. 5).
3.2 Material Model
The medium $\Omega$ is modeled as incompressible neo-Hookean. Strain energy density for an incompressible neo-Hookean material is
$W=\frac{\mu}{2}(I_{1}-3),$ (2)
where $\mu$ is the shear modulus of the material, $I_{1}=\text{tr}(\textbf{B})$ is the first invariant of the left Cauchy-Green deformation tensor $\textbf{B}=\textbf{F}\textbf{F}^{T}$, and F is the deformation gradient tensor with $\det\textbf{F}=1.$
The analytical dimensionless pressure-volume relation for the growth of a spherical cavity in an infinite incompressible neo-Hookean medium is given as [13, 41]
$\overline{p}=\frac{1}{2}(1-\overline{V}^{-4/3})+2(1-\overline{V}^{-1/3}).$ (3)
where it is seen that the bulk cavitation limit can be found by setting $\overline{V}\to\infty$ as $\overline{p}_{\rm c,bulk}=2.5$.
4 Computational Framework
4.1 Discretization and Simulation Setup
We carry out the simulations using the open source finite element software FEniCS [1, 30, 31]. Incompressibility is enforced via a mixed displacement-pressure ($u-\tilde{p}$) formulation using the lowest-order Taylor-Hood element pair (quadratic elements for displacement and linear elements for pressure), with a perturbed Lagrangian treatment of the volumetric constraint ($J-1=\tilde{p}/K$, where $\tilde{p}$ is the Lagrange multiplier interpreted as a hydrostatic pressure and $K=10^{12}\times G$ is a penalty parameter with the interpretation of a bulk modulus). The weakform setup in FEniCS parallels the implementation in [42] (see also [2]). The domain dimensions used in the simulation are shown in Figure 5. The mesh is especially refined near the cavity to ensure sufficient accuracy in capturing the large deformations in the vicinity of the cavity. Figure 6 shows an example of the meshes employed; here, the minimum element size is $R/250$, and 168758 triangular elements are used to discretize the domain using the software Gmsh [15]. Mesh convergence study is presented in A.
Figure 6: Spatially adaptive mesh discretization ($d/R=7.5$).
The boundary value problems illustrated in Figure 5 are numerically solved for both rigid and symmetric interfaces for dimensionless distance values $d/R\in\{$1.10, 1.25, 1.41, 1.60, 1.82, 2.06, 2.33, 2.65, 3.00, 3.37, 3.78, 4.24, 4.76, 5.35, 6.00, 6.63, 7.33, 8.11, 8.96, 9.91, 10.95, 12.11, 13.39, 14.80, 16.36, 18.09, 20.00}\}. The pressure–volume data are then extracted (see A), resulting in the plots shown in Fig. 2 and Fig. 3.
4.2 Cavitation Pressure through Curve Extrapolation
Given that the simulations show a monotonic increase in cavity volume with applied pressure, where the pressure appears to asymptote in all cases, the cavitation pressure can be determined by the limiting value of the pressure as $\overline{V}\to\infty$. While the simulations are able to achieve large values of $\overline{V}$, the pressure-volume curves need to be extrapolated to determine the asymptotic cavitation pressure value. We achieve this by employing a curve-fitting procedure for the pressure-volume curve and then using the fitted curve to estimate the cavitation pressure.
Motivated by the expression for the pressure-volume relation for the bulk cavitation case in eq. (3), we employ polynomial basis functions in terms of the variable $\overline{V}^{-1/3}$ for the fitting. Specifically, we choose Laguerre polynomials of degree $m$, given by
$L_{m}(x)=\sum_{k=0}^{m}\binom{m}{k}\frac{(-1)^{k}}{k!}x^{k},$ (4)
since they are orthonormal in $x\in[0,\infty)$, where the functional fit for the dimensionless pressure-volume curve is given as
$\hat{p}(\overline{V})=\sum_{m=0}^{N}a_{m}L_{m}(\overline{V}^{-1/3}),$ (5)
and $a_{m}$ are the fitting coefficients. Given that $L_{m}(0)=1$, the cavitation pressure is determined by the sum of the coefficients as
$\overline{p}_{\rm c}=\lim_{\overline{V}\to\infty}\hat{p}=\sum_{m=0}^{N}a_{m}$ (6)
The root-mean-square error (RMSE) of the fitted curve is defined as
$\mathrm{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\overline{p}_{i}-\hat{p}_{i}\right)^{2}}$ (7)
where $\hat{p}_{i}\coloneqq\hat{p}(\overline{V}_{i})$ is the pressure predicted by the fitted model, $i$ denotes the index of the data points $(\overline{p}_{i},\overline{V}_{i})$ in the discrete numerical dataset, and $n$ is the number of data points in the dataset.
We find that the RMSE decreases with increasing number of fitting terms $N+1$, however for $N>5$, the condition number of the matrix associated with the linear equation system to solve for the least squares fitting of the parameters $a_{m}$ becomes poor, reflecting over-parameterization and poor numerical stability of the fitted coefficients. Consequently, we cap the value of $N$ at 5, which gives satisfactory fit (low RMSE) for all $d/R$ values and report the average of the cavitation pressures obtained using the fits for $N=4$ and $N=5$ (the choice of $N\geq 4$ covers all the terms in eq. (3)). Further, we note that using a regular polynomial basis instead of the Laguerre polynomial basis yields identical results reflecting the robustness of the fitted cavitation pressure. The fitting procedure is illustrated for two different $d/R$ values, for the cavity-interface interaction problem in Fig. 7-7, and for the cavity-cavity interaction problem in Fig. 7-7.
[…]
Figure 7: Fitting procedure results for: (a) cavity-interface, $d/R=20$ (b) cavity-interface, $d/R=1.1$ (c) cavity-cavity, $d/R=20$ and (d) cavity-cavity, $d/R=1.1$.
5 Conclusion
We consider two extensions of the classical cavitation problem to analyse the effects of a rigid interface and, separately, another identical cavity on the critical pressure for cavitation in an incompressible, infinite neo-Hookean medium. Through axisymmetric finite element simulations and a curve-fitting based (Laguerre polynomials basis functions) extrapolation procedure, we calculate the cavitation pressure as a function of the cavity-interface distance and inter-cavity half-distance, respectively.
For a cavity near a rigid interface, the cavitation pressure decreases monotonically with increasing cavity–interface spacing, with values ranging between $p_{\rm c}\approx 3.44\mu$ for small d/R(=1.1)d/R~(=1.1) and bulk cavitation value of $p_{\rm c,bulk}=2.5\mu$ for large $d/R$. In cavity-cavity interaction case, the cavitation pressure is non-monotonic as a function of inter-cavity half-distance, with a peak cavitation pressure $p_{\rm c}\approx 2.67\mu$ exhibited around $d/R\approx 3.8$. The noticeable enhancement effect in both cases, with respect to the classical bulk cavitation pressure, indicates the importance of taking into account the presence of interface and nearby cavities, which is in turn likely to affect the mechanical response and failure in a variety of engineering and biological systems.
Appendix A Mesh Sensitivity Analysis
Four levels of mesh refinement were tested for a representative value of $d/R=7.5$ as shown in Figure 8.
Figure 8: Different levels of adaptive refinement for $d/R=7.5$: (a) coarse (b) medium (c) fine (d) finest.
The remote pressure at the top boundary is calculated as
$p=p_{\mathrm{remote},\mathrm{ave}}=\frac{\int_{X_{2}=L+d}\,(\bm{\sigma}\mathbf{n}.\mathbf{n})\,\mathrm{d}a(\mathbf{X})}{\int_{X_{2}=L+d}\,\mathrm{d}a(\mathbf{X})}$ (8)
where $\bm{\sigma}$ is the Cauchy stress tensor, $\mathbf{n}$ is the outward unit normal in the deformed configuration, $\mathbf{X}$ is the position vector of material points in the undeformed reference configuration, and $\mathrm{d}a$ is a deformed area of a differential surface element which is defined as
$\mathrm{d}a=2\pi r_{0}\,J\|\bm{\mathrm{F}}^{-T}\bm{\mathrm{N}}\|\,\mathrm{d}S$ (9)
where $r_{0}$ is the undeformed coordinate in $X_{1}$ direction, $\bm{\mathrm{F}}$ is the deformation gradient with $J=\mathrm{det}\bm{\mathrm{F}}(\approx 1)$, the outward unit normal in the reference configuration is $\bm{\mathrm{N}}$, and $\mathrm{d}S$ is the length of the differential arc element in the 2D axisymmetric representation. Moreover, the deformed volume of the cavity is determined as
$V=\int_{X_{1}^{2}+(X_{2}-d)^{2}=R^{2}}\pi r^{2}\left(-n_{1}\right)\,\mathrm{d}s(\mathbf{X})$ (10)
where $r$ is the deformed coordinate in $X_{1}$ direction, $n_{1}$ is the component of the outward unit normal $\bm{\mathrm{n}}$ in $X_{1}$ direction, and $\mathrm{d}s$ is the deformed length of the differential arc element in the 2D axisymmetric representation. The deformed and undeformed arc lengths in the 2D axisymmetric plane are related through the relation
$\bm{\mathrm{n}}\,\mathrm{d}s=\textrm{det}(\bm{\mathrm{F}}_{\rm 2D})\bm{\mathrm{F}}_{\rm 2D}^{-T}\bm{\mathrm{N}}\,\mathrm{d}S.$ (11)
where $\bm{\mathrm{F}}_{\rm 2D}$ is the deformation gradient in the 2D axisymmetric plane related to the full deformation gradient as follows,
$\bm{\mathrm{F}}=\bm{\mathrm{F}}_{\rm 2D}+\frac{r}{r_{0}}\hat{e}_{3}\otimes\hat{e}_{3}$ (12)
where $\hat{e}_{3}$ is a unit vector along $X_{3}$ axis.
The dimensionless pressure vs volume curves obtained for the different levels of mesh refinement are shown in Fig. 9 for the cavity-cavity interaction case. In addition, the asymptotic cavitation pressure is determined and shown in the figure using the corresponding colours. There is minimal change in the critical pressure estimate with mesh refinement and the remote pressure is well estimated even with a ‘coarse’ mesh.
[…]
Figure 9: Cavity-cavity interaction problem for $d/R=7.5$ and different levels of mesh refinement: (a) dimensionless pressure-volume curves, and (b) average dimensionless cavity pressure against dimensionless volume.
Next we check the traction free boundary condition at the cavity surface. The average pressure in the cavity is calculated as
$p_{\mathrm{cavity},\mathrm{ave}}=\frac{\int_{X_{1}^{2}+(X_{2}-d)^{2}=R^{2}}(\bm{\sigma}\mathbf{n}.\mathbf{n})\,\mathrm{d}a(\mathbf{X})}{\int_{X_{1}^{2}+(X_{2}-d)^{2}=R^{2}}\mathrm{d}a(\mathbf{X})}$ (13)
The average cavity pressure is presented in Fig. 9 as a function of dimensionless deformed cavity volume, for the different levels of mesh refinement. The “finest” mesh provides an acceptable accuracy in satisfying the natural boundary condition at the cavity, with an error of less than one percent. All results shown in the manuscript are using the finest level of mesh refinement.